Make a use of calculus to prove that the price elasticity of demand is a constant ε everywhere along the demand curve whose function is Q = Apε , where A is a positive constant and p is the market price


1
Expert's answer
2021-07-20T10:47:28-0400

Differentiating the demand function as Q = Apε with respect to p, we find that "\frac{dQ}{dp} =ε Apε^{-1}" . To get the elasticity, we multiply "\frac{dQ}{dp}" by "\frac{p}{Q} = \frac{p}{Apε^{-1}} =\frac{1}{Apε^{-1}}" That is, the elasticity is

"ε Apε^{-1} \times \frac{1}{Apε^{-1}} =ε ." Because this result holds for any p, the elasticity is the same, ε, at every point along the demand curve.